OpenStudy (anonymous):

Evaluate the expression when x = 8 and y = 1.

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

err.... which mean you just need to replace the x and y to the equations by it numbers like: give x=2 and y =7 7x+ 3y =5y 7(2) +3(7) = 5(7) 14+21=35 35=35 or something similar like that